# A Strong Baseline for Temporal Video-Text Alignment

### v1

@@ 5 Conclusion
To conclude, we have established a simple Transformer-based architecture for video-text alignment in large-scale instructional videos. We have investigated the factors that can potentially affect the performance, for example, upgrading the ASR system, exploring various visual-language backbones, including CLIP, S3D, InterVideo, transforming the noisy ASR transcript into descriptive steps with post-determined temporal windows as extra training source. When evaluating on three different public benchmarks, our method surpassing the existing state-of-the-art methods by a significant margin.
Appendix
Appendix A Additional Implementation Details
A.1 Architecture
@@
In this section, we provide more details on the visual-language backbones, and the alignment module, as introduced in Sec. 3.3 (main paper).
@@
Visual Backbone. We adopt S3D-G pre-trained with MIL-NCE, CLIP ViT/L-14 pre-trained with InfoNCE and InternVideo-MM-L-14 pre-trained with reconstruction and contrastive loss. For S3D-G, the original video is first resized to have a shorter side equal to 256 pixels, and then the frames are center-cropped with resolution before feeding into the S3D-G. For CLIP and InternVideo, the original video is first center cropped to a shorter side and then resized to resolution.
@@
Textual Backbone. The text encoder associated with S3D-G adopts a Bag-of-word (BoW) model based on Word2Vec embeddings pre-trained on GoogleNews. Each sentence is tokenized and truncated under 32 tokens. After the embedding layer, each word vector is further projected by fully-connected layers, and a max-pooling layer is applied to get one sentence embedding. For the text encoder of CLIP and InternVideo, it takes a maximum of 77 tokens for each sentence. Each token is passed to an embedding layer and added with positional encoding. After being encoded by the Transformer blocks, we take sentence features from the <eot> embedding for each sequence.
@@
Alignment Modules. For video-text alignment modules in our proposed model, we use 3-layer Transformer encoder blocks and 3-layer decoder blocks with 8-head attention. The model dimension equals 256, and the projection dimension to compute the cosine similarity equals 64. At training time, we will truncate the videos whose duration is longer than 1200 seconds and set the batch size as 8. Our model is trained on a single GPU (NVIDIA GeForce RTX 3090) for approximately 20 hours on the full training data (WhisperX + HowToStep).
A.2 Dataset Details
HowToStep (Training). For our final HowToStep, the average steps (sentences) per video is 10.6 and the average words per step is 8.0. As shown in Fig. 4(a) and Fig. 4(b), we plot the distribution of steps per video and words per step. In addition, we present a word cloud to show the descriptions of the summarized steps in HowToStep (Fig. 4(c)).
@@ A.2 Dataset Details
HTM-Align (Evaluation). Given an instructional video from HowTo100M with its narrations with start-end timestamps from the Youtube ASR transcript, the annotator determines whether each narration is alignable with the video (i.e., visually related) and adjusts the ground truth temporal window to cover the visual content if the narration is alignable. The metric is Recall@1, which evaluates whether the predicted timestamp of the narrations that are alignable with the video falls into the ground truth temporal window [han2022temporal].
@@
HT-Step (Evaluation). Given the video from HowTo100M with the task name (e.g., Make Pumpkin Puree) and the recipe steps from the corresponding Wikihow article, the annotator decides whether the video is relevant to the task. If the video is relevant to the task of steps, the annotator will mark all the instances of the steps with a temporal window. The test set contains 600 videos, with 5 videos per task and the metric is Recall@1 as well. The complete statistics for this dataset can be found in [mavroudi2023learning].
@@
CrossTask (Evaluation). The CrossTask dataset consists of 4800 videos from 18 primary activities and 65 related activities. The videos of primary activities are annotated with steps from a predefined taxonomy of 133 atomic steps (e.g., add onion, add taco) and the corresponding temporal windows in the video. For the step localization task on this dataset, the metric is Average Recall@1, which means computing Recall@1 for steps of videos from each primary activity and averaging the results of different activities. Following previous work [mavroudi2023learning], we report the average results over 20 random sets of 1850 videos from 18 primary activities.
A.3 Task Details
@@
For the narration alignment task, the texts are in the form of narrations in videos, with strong temporal correlations. The consecutive sentences in narrations typically follow a temporal order. However, in procedural step grounding, the texts are in the form of steps collected from a knowledge base, namely, WikiHow. Due to the discrepancy between the knowledge base and the instructional videos, the order of steps in the knowledge base may differ from the order of actions in the videos. Therefore, the temporal order of consecutive sentences is not fixed. Due to the different temporal attributes of the texts in the two tasks, we employ distinct training techniques for each task. As described in Sec. 3.3.2, during training for the narration alignment task, we add learnable positional embeddings to textual features and input all the texts of one video into the alignment module with a temporal order, which is determined differently for WhisperX texts (based on the ASR recognition time order) and HowToStep texts (based on temporal windows determined in the 2-stage temporal grounding). However, for the procedural step grounding task, we do not add positional embeddings, and input the texts of one video into the alignment module after shuffling their sequential order in training.
Appendix B Extra Ablations
Ablation on prompts for LLM on dataset construction. Due to the excessive resources required for conducting prompt ablations on the complete HTM-370K dataset, which contains 370k videos, we perform the prompt ablations on approximately 16% of the total 370k videos, amounting to 61k videos. We use the same hyper-parameters in the 2-stage temporal grounding procedure for generated steps and then train on the same model to get the results on HT-Step. In Tab. 6, we experiment with various prompts to guide the LLM in generating descriptive steps for ablation. The first row in Tab. 6 represents the basic prompt, which introduces the background of the task, a brief description of the task, and the general format requirements for output. Subsequently, by incorporating key phrases like “each step should be a short and concise phrase”, “Do not output colloquial sentences in the speech” and “The number of steps in your output should range from zero to four steps”, the LLM is guided to distill concise steps from the redundant ASR transcripts, thereby leading to performance improvement. Hence, we ultimately choose the prompt in the fourth row of Tab. 6 to generate descriptive steps.
Ablation on temporal grounding. As mentioned in Sec. 3.2.2, we propose a two-stage method to determine the start/end timestamp for each generated descriptive step. Here, we conduct more ablation studies on the filtering thresholds used for temporal grounding, i.e. and . We choose the best thresholds ( and ) to generate the final dataset. In Tab. 7 we ablate the complete design choices for the two-stage temporal grounding approach, including the filtering threshold, the position of maximal alignment score in each step segment, and the duration of time allocated to each step. We can observe that setting the maximum similarity as the starting timestamp of the step with a constant duration brings the best performance.
Prompt |
HT-Step |
| R@1 | |
@@ Appendix B Extra Ablations
: I will give you an automatically recognized speech with timestamps from a video segment that is cut from a long video. : The speaker in the video is teaching the audience to do something. : Your task is to summarize the key steps in order. : Output only the numbered key steps without timestamps. : Here is this automatically recognized speech: ASR transcript |
38.8 |
: Each step should be short and concise phrase. |
38.8 |
: Do not output colloquial sentences in the speech. |
39.5 |
: The number of steps of your output should be between zero step and four steps. |
40.3 |
@@
| HT-Step R@1 | 36.0 | 35.7 | 37.6 | 42.6 | 43.0 | 44.0 | 35.8 | 35.6 | 36.6 | |
Appendix C Qualitative Results
C.1 ASR Transformation Example
In Fig. 6, we provide an example of using an LLM with the prompt to transform ASR transcripts into descriptive steps. Comparing ASR transcripts with descriptive steps, we can observe that even ASR transcripts with completely accurate recognition can have issues of semantic ambiguity and redundancy. For instance, in Fig. 2a, sentences in orange are irrelevant to the task at hand, while those highlighted in blue exhibit issues with unclear references. Additionally, a segment of speech recognized by the ASR system often contains many important actions corresponding to different time intervals in the video, which is a significant source of misalignment between text and video. Conversely, the descriptive steps generated by the LLM are more concise and clear, eliminating the redundant information in the ASR transcripts and articulating the action procedure a unified form.
C.2 Alignment Visualization
@@
Appendix D Limitations and Ethical Concerns
As a proposal-free method for video-text alignment, we do not explicitly generate the temporal window for each narration or procedural step, and only obtain it via post-processing the alignment score matrix. In addition, we pay more attention to whether the most possible timestamp for each text falls into the ground truth temporal window since the start-end boundary of the text is not used under the Recall@1 metric compared with Recall@IoU metrics. For ethical concerns, we are aware that the public instructional video dataset and the knowledge of large language models may have gender, age, geographical, or cultural bias.

### v2

@@ 5 Conclusion
To conclude, we have established an automatic pipeline for constructing a high-quality video-text dataset for multi-sentence grounding in large-scale instructional videos. We have investigated the factors potentially affecting performance, including upgrading the ASR system, transforming the noisy ASR transcripts into descriptive steps by LLMs as an extra training source, and proposing a simple Transformer-based model to refine the temporal windows for each step. When evaluating three public benchmarks of multi-sentence grounding tasks, our method surpasses the existing state-of-the-art methods by a significant margin.
Acknowledgements
This work is supported by National Key R&D Program of China (No.2022ZD0161400).
Supplementary Materials
6 Additional Implementation Details
6.1 Architecture Details
@@
In this section, we provide more details on the visual-language backbones, and the grounding module, as introduced in Sec. 3.2 (main paper).
@@
Visual Backbone. We adopt S3D-G pre-trained with MIL-NCE, CLIP ViT/L-14 pre-trained with InfoNCE and InternVideo-MM-L-14 pre-trained with reconstruction and contrastive loss. For S3D-G, the procedure begins by decoding the original video at 16 frames per second (fps), followed by resizing it to ensure the shorter dimension is 256 pixels. After resizing, the frames are then center-cropped to a resolution of before being inputted into the S3D-G. Each 16-frame video clip cropped by a non-overlapping temporal window is fed into the S3D-G architecture, resulting in one feature (512-d) per second. For CLIP and InternVideo, the original video is first center cropped to a shorter side and then resized to resolution. When using CLIP, we decode the video into 1fps and extract one visual feature (768-d) per second with OpenAI’s ViT/L-14 model [13]. As for InternVideo, We decode the video into 8fps and feed it into the InternVideo-MM-L14 model with an 8-frame non-overlapping temporal window, obtaining one visual feature (768-d) per second.
@@
Textual Backbone. The text encoder associated with S3D-G adopts a Bag-of-word (BoW) model based on Word2Vec embeddings pre-trained on GoogleNews. Each sentence is tokenized, truncated under 32 tokens, and then encoded with the text encoder associated with the S3D-G video encoder [28]. Specifically, the sentence is tokenized and then converted into word vectors (300-d) through an embedding layer trained on Google News. Following this, word vectors from the same sentence will be projected into 512-d vectors, and turned into one sentence vector through maxpooling. For the text encoder of CLIP and InternVideo, it takes a maximum of 77 tokens for each sentence. Each token is passed to an embedding layer and added with positional encoding. After being encoded by the Transformer blocks, we take sentence features from the <eot> embedding for each sequence.
@@
Grounding Modules. For multi-sentence grounding modules in our proposed model, we use 3-layer Transformer encoder blocks and 3-layer decoder blocks with 8-head attention mechanisms. The model dimension, denoted as , is set to 256, while the projection dimension, represented as , utilized for computing the cosine similarity, is 64. At training time, we will truncate the video, whose duration is longer than 1200 seconds, and set the batch size as 8. Our model is trained on a single GPU (NVIDIA GeForce RTX 3090) for approximately 20 hours on the whole training data (WhisperX + HowToStep).
6.2 Dataset Details
HowToStep (Training). In our final HowToStep, the average number of steps (sentences) per video stands at 10.6, with an average word count per step amounting to 8.0. As shown in Fig. 5(a) and Fig. 5(b), we plot the distribution of steps per video and words per step. In addition, we present a word cloud to show the descriptions of the summarized steps in HowToStep (Fig. 5(c)). To validate the diversity of procedural steps in our dataset, we conduct a quantitative comparison with the latest work [2] in Fig. 6(a), by comparing the number of aligned steps within the same set of videos. On five tasks that have the most Wikihow steps, our step count is 11.2 times more than [2] on average. Additionally, using t-SNE to reduce sentence embedding dimensions for the ’How To Make Fried Pickles’ task, our steps are semantically more diverse, as shown in Fig. 6(b).
@@ 6.2 Dataset Details
HTM-Align (Evaluation). Given an instructional video from HowTo100M with its narrations with start-end timestamps from the YouTube ASR transcript, the annotator determines whether each narration is alignable with the video (i.e., visually related) and adjusts the ground truth temporal window to cover the visual content if the narration is alignable. We employ the metric, Recall@1, to evaluate whether the predicted timestamp of the narrations are alignable with the video falls into the ground truth temporal window [17].
@@
HT-Step (Evaluation). Given the video from HowTo100M with the task name (e.g., Make Pumpkin Puree) and the recipe steps from the corresponding Wikihow article, the annotator decides whether the video is relevant to the task. If the video is relevant to the task of steps, the annotator will mark all the instances of the steps with a temporal window. The test set contains 600 videos, with 5 videos per task and the metric is Recall@1 as well. The complete statistics for this dataset can be found in [27].
@@
CrossTask (Evaluation). The CrossTask dataset consists of 4800 videos from 18 primary activities and 65 related activities. The videos of primary activities are annotated with steps from a predefined taxonomy of 133 atomic steps (e.g., add onion, add taco) and the corresponding temporal windows in the video. For the step localization task on this dataset, the metric is Average Recall@1, which means computing Recall@1 for steps of videos from each primary activity and averaging the results of different activities. Following previous work [27], we report the average results over 20 random sets of 1850 videos from 18 primary activities.
6.3 Task Details
@@
For the narration alignment task, the texts are in the form of narrations in videos, with strong temporal correlations. The consecutive sentences in narrations typically follow a temporal order. However, in procedural step grounding task, the texts are in the form of steps collected from a knowledge base, namely, WikiHow. Due to the discrepancy between the knowledge base and the instructional videos, the order of steps in the knowledge base may differ from the order of actions in the videos. Therefore, the temporal order of consecutive sentences is not fixed. Due to the different temporal attributes of the texts in the two tasks, as described in Sec. 3.2, we have delineated an indicator to ascertain the inclusion of learnable positional encodings for textual features. Specifically, we have defined a token that denotes the type of text. When the token is denoted as <narration>, signifying that the input bears pronounced temporal correlations akin to a narration, we integrate positional encodings and, during training, input all the texts of one video into the grounding module in their temporal order. Conversely, when the token is represented as <step>, indicating that the input text does not follow a temporal sequence, we refrain from incorporating positional encodings and input the texts of a single video into the grounding module in a shuffled sequence for training purposes. Throughout the training phase, we randomly allocate the type-token of the training samples as <narration> with a 50% probability, and as <step> with the remaining 50% probability. During the inference stage of the narration alignment task, we classify the type-token for all samples as <narration>; conversely, in the inference stage of the procedural step grounding task, we assign the type-token for all samples as <step>. This methodology allows us to achieve both tasks with a unified set of weights.
Prompt | Two-Stage | HT-Step |
Determination | R@1 | |
I will give you an automatically recognized speech with timestamps from a video segment that is cut from a long video. Write a summary for this video segment. Write only short sentences. Describe only one action per sentence. Keep only actions that happen in the present time. Begin each sentence with an estimated timestamp. Here is this automatically recognized speech: timestamp + ASR transcript | ✗ | 19.6 |
@@
I will give you an automatically recognized speech with timestamps from a video segment that is cut from a long video. The speaker in the video is teaching the audience to do something. Your task is to summarize the key steps in order. Each step should be short and concise phrase. Do not output colloquial sentences in the speech. Output only the numbered key steps without timestamps. Here is this automatically recognized speech: timestamp + ASR transcript | ✓ | 41.8 |
I will give you an automatically recognized speech from a video segment that is cut from a long video. The speaker in the video is teaching the audience to do something. Your task is to summarize the key steps in order. Each step should be short and concise phrase. Do not output colloquial sentences in the speech. Describe only one action per sentence. Output the numbered key steps. Here is this automatically recognized speech: ASR transcript | ✓ | 43.7 |
7 Extra Ablations
Ablation on dataset construction. In this section, we perform the ablations on HTM-370K, concerning the prompt for guiding the LLM in generating descriptive steps and the methodology for equipping the summarised procedural steps with corresponding video timestamps. A straightforward method to acquire descriptive steps and their corresponding video segments is to feed both ASR transcripts and their respective timestamps into a LLM, prompting the LLM to simultaneously generate descriptive steps along with their associated timestamps. Specifically, we utilize the format timestamp + ASR transcript as input, employing the prompt in the first row of Tab. 6 to guide the LLM in generating descriptive steps, along with timestamp for each step. However, given that timestamps from ASR transcripts may not be aligned, and LLM cannot rectify this misalignment based solely on text, the steps generated by this method cannot endow the model with proficient multi-sentence grounding capabilities. An alternative approach involves the LLM solely undertaking the task of step generation, followed by determining the timestamp for each step through the two-stage determination procedure proposed in Sec. 3.1 and Sec. 3.3. Employing this approach, we devised two prompts, as delineated in the second and third rows of Tab. 6. The distinction lies in that for the second row, we provide the corresponding timestamp extracted from the ASR, with the intention of guiding the LLM to utilize temporal information for generating the steps; whereas, in the third row, we omit any input of timestamp information. We use the same hyper-parameters in the 2-stage temporal grounding procedure for generated steps and then train on the same model to get the results on HT-Step. From Tab. 6, it is evident that steps generated by the LLM without incorporating timestamps result in improved grounding capabilities for the model. This could be attributed to inherent misalignment within the timestamps extracted from ASR transcripts, which may inadvertently misguide the inference process of the LLM. On the other hand, when timestamps are included in the input, the outcomes generated by the LLM frequently contain timestamp information, despite restrictions imposed in the prompts against this. Hence, we ultimately choose the prompt in the third row of Tab. 6 to generate descriptive steps.
Ablation on LLM choice in proposed pipeline. In this section, we performed an ablation study on roughly 10% of the total training dataset, employing different LLMs, namely Llama-2, Mistral, and Llama-3. As shown in Tab. 7, Mistral and Llama-2 demonstrated similar performance, while Llama-3 exhibited the best performance. We adopt Llama-2, but the quality of data generated by our proposed pipeline can continuously improve with the advancement of LLMs.
LLM | HT-Step R@1 | HTM-Align R@1 |
Llama-2 | 35.5 | 34.9 |
Mistral | 35.1 | 34.7 |
Llama-3 | 35.7 | 36.7 |
Ablation on temporal grounding. As mentioned in Sec. 3.1 and Sec. 3.3, we propose a two-stage method to determine the start/end timestamp for each generated descriptive step and we can repeat the refined process (i.e., self-training) multiple iterations. In Tab. 8 we ablate the complete design choices for the first iteration of temporal refinement, including the filtering threshold, the position of maximal alignment score in each step segment, the duration of time allocated to each step. Setting the maximum similarity as the starting timestamp of the step with a constant duration brings the best performance for the first interation refinement. Subsequently, we conduct ablation study on the iterations of self-traininig, with each refinement round considering various filtering thresholds, positions of maximal alignment score, and step durations as what we do in Table 8. Consequently, we present the best result for each iteration in Tab. 9. It is evident that performing two iterations of temporal refinement yields the best performance. However, it is worth noting that in all the aforementioned experiments, we chose only one round as the default to balance computation cost and performance improvement.
@@ 7 Extra Ablations
HT-Step R@1 | 37.6 | 37.5 | 38.2 | 43.1 | 43.7 | 43.0 | 38.0 | 37.2 | 37.8 |
Iters | thresh | pos | Training | HT-Step | HTM-Align | |
Text | R@1 | R@1 | ||||
✗ | 0.2 | center | ✗ | S | 36.0 | 44.0 |
1 | 0.8 | start | 8 | S | 43.7 | 46.7 |
1 | 0.8 | start | 8 | W + S | 46.4 | 71.6 |
2 | 0.7 | center | ✗ | S | 44.8 | 52.9 |
2 | 0.7 | center | ✗ | W + S | 46.9 | 72.1 |
3 | 0.8 | center | ✗ | S | 45.2 | 51.8 |
3 | 0.8 | center | ✗ | W + S | 46.0 | 69.7 |
8 Qualitative Results
8.1 ASR Transformation Example
In Fig. 7, we provide an example of using a LLM with the prompt to transform ASR transcripts into descriptive steps. Compared with descriptive steps, some complete accurately recognised ASR transcripts still contain semantic ambiguity and redundancy. For instance, in Fig. 2a, sentences in orange are irrelevant to the task at hand, while those highlighted in blue exhibit issues with unclear references. Additionally, a segment of speech recognized by the ASR system often contains many important actions corresponding to different time intervals in the video, which significantly cause misalignment between text and video. Conversely, the descriptive steps generated by the LLM are more concise and clear, eliminating the redundant information in the ASR transcripts and articulating the action procedure in a unified form.
8.2 Alignment Visualization
@@
9 Limitations and Ethical Concerns
As a proposal-free method for multi-sentence grounding, we do not explicitly generate the temporal window for each narration or procedural step, and only obtain it via post-processing the alignment score matrix. In addition, we pay more attention to whether the most possible timestamp for each text falls into the ground truth temporal window since the start-end boundary of the text is not used under the Recall@1 metric compared with Recall@IoU metrics. For ethical concerns, we are aware that the public instructional video dataset and the knowledge of large language models may have gender, age, geographical, or cultural bias.